# A Comprehensive Survey in LLM(-Agent) Full Stack Safety: Data, Training and Deployment

### v1

@@ 6.1.2 Defensive Mechanisms in Deployment
Output Filtering Mechanisms Output filtering mechanisms [193, 613] serve as a critical safeguard in LLM deployment, ensuring that generated responses comply with safety constraints while preserving informativeness. Unlike input preprocessing, which aims to prevent adversarial prompts from reaching the model, output filtering mitigates harmful content post-generation. Existing approaches primarily follow three paradigms: rule-based constraints, generative adversarial filtering, and toxicity detection.
@@
Toxicity detection [626, 627, 628] and content moderation [629, 630, 631] further reinforce output safety by identifying and mitigating hate speech [632], misinformation, and other harmful content. Supervised fine-tuning adapts LLMs to recognize undesirable patterns, while classifier-based detection models [633] filter responses in real-time. Some approaches introduce debiasing strategies, such as controlled decoding [634, 635] and anti-expert guidance [636], to suppress toxic outputs without sacrificing response diversity. However, these methods face challenges in balancing false positives and false negatives, particularly in ambiguous or context-dependent cases.
@@
Robust Prompt Engineering Robust prompt engineering aims to enhance LLM safety by designing input prompts that resist adversarial manipulation [637], protect sensitive data, and mitigate harmful outputs—all [638] without modifying model parameters. These strategies act at the interaction level, offering lightweight and model-agnostic protection.
@@
Privacy-preserving prompt [648] design introduces formal guarantees through differential privacy. Approaches like DP-Prompt [649] and stochastic gradient masking [650] reduce information leakage from prompts without harming performance. Desensitization and directional control of in-context representations offer additional privacy protections during prompt construction. Prompt engineering [534] also helps mitigate societal risks. Chain-of-thought prompting and guided templates reduce gender bias [651] in reasoning tasks, while prompt learning [652] improves toxicity detection and generation control [653, 654], often surpassing specialized models in efficiency and generalization.
@@
System-level Security Controls System-level defenses [659] enhance LLM deployment by optimizing inference, enforcing alignment, isolating untrusted inputs, and securing the supply chain. Systems like Petals [660], Sarathi-Serve [661], and DistServe [662] restructure computation to improve throughput and latency, while TriForce [663], Medusa [664] MagicDec [665] accelerate generation via speculative decoding and structural compression. Parallel frameworks such as DeepSpeed-FastGen [666] and SpecExec [667] further boost efficiency with minimal overhead.
@@ 6.1.3 Evaluation and Benchmarks in Deployment
To systematically assess the reliability of large language models (LLMs) after deployment, we categorize robustness evaluation into two broad types: adversarial robustness and natural robustness. Adversarial robustness focuses on evaluating how LLMs respond to malicious or adversarial inputs, such as jailbreak prompts, prompt injections, or red-teaming attacks. Natural robustness, on the other hand, assesses LLM behavior under non-malicious but realistic distribution shifts, including typos, paraphrasing, novel word usage, or temporal drift. A summary of representative benchmarks categorized along these two dimensions is presented in Table VII.
@@
Content Trustfulness and Fairness Evaluation
@@
Data Privacy and Leakage Evaluation
@@
Multi-modal Safety Evaluations As multimodal large language models (MLLMs) become increasingly integrated into real-world applications, ensuring their safety under diverse input conditions is essential. A growing number of studies have proposed evaluation benchmarks and frameworks to assess MLLM vulnerabilities across multiple dimensions [708, 709, 710, 711, 712, 713, 714, 715, 716, 717, 718, 719, 720, 721, 722, 723, 724, 725, 726, 727, 728].
@@ 6.2 Single-agent Safety
In this section, we focus on security issues related to a single agent. We first define an agent as an interactive entity that uses an LLM as the core for reasoning, decision-making, and reflection while integrating memory, tools, and the environment as capability-enhancing components. Beyond the deployment risks associated with the LLM core, we introduce the security issues arising from these three additional modules. Specifically, for tools (Section 6.2.2) and memory (Section 6.2.3), we summarize existing work from both attack and defense perspectives to identify technical paradigms. For the environment (Section 6.2.6), we explore unique security challenges from the perspective of various agent-interaction settings. We demonstrate an overview of agent safety in Figure 12.
@@ 6.2.2 Tool Safety
Attacks Based on the target of the attack, safety-related attacks involving tools can be categorized into Tool-aided Attacks and Tool-targeted Attacks. The former refers to attackers utilizing agents equipped with tools to execute attacks that LLMs cannot independently assist with, such as leveraging agents with web access and code execution capabilities to facilitate cyberattacks. The latter involves attackers targeting the tool invocation process itself, attempting to manipulate or induce tool selection for malicious purposes through various attack methods. However, from the perspective of the technical stack of attacks, the two can be unified. We have identified new applications of traditional LLM attack methods in tool safety, as well as novel attack paradigms that have emerged due to the unique characteristics of tools.
@@
Defenses Compared to attacks on agent tools, defense mechanisms for secure tool invocation have been less studied. Specifically, AgentGuard [766] employs LLM orchestrators to automatically detect unsafe tool-use workflows and produce safety constraints for secure tool utilization. PrivacyAsst [767] proposes an encryption-based solution by integrating an encryption scheme into the tool using LLM agents to safeguard user privacy and align them with computational security standards. In addition, some works enhance the security of agent systems by leveraging tool invocation, GuardAgent [768] pioneers an approach to verify target agents’ trustworthiness by executing guardrail code through API calls during task plan implementation.
@@ 6.2.3 Memory Safety
(I) Memory Poisoning refers to adversarial attacks where malicious data is injected into an agent’s long-term memory [771, 772, 773, 774, 284, 775]. When the agent retrieves and utilizes such corrupted memory, it may produce erroneous outputs, misleading responses, or even hazardous actions. For example, the PoisonedRAG framework [773] employs a dual optimization approach, simultaneously manipulating both the retrieval and generation pipelines to systematically poison the agent’s memory system. AgentPoison [772] introduces an advanced backdoor attack methodology that optimizes trigger patterns and seamlessly integrates them into query formulations, significantly elevating the likelihood of malicious sample retrieval while maintaining stealth. (II) Privacy Leakage occurs when attackers exploit the interface between an agent and its long-term memory to extract stored sensitive data [776, 562, 777, 479, 560]. Such breaches may expose user information to malicious third parties, posing significant real-world risks. (II) Memory Misuse refers to the deliberate construction of multi-turn query sequences that systematically circumvent safety protocols by exploiting the retention properties of agent short-term memory [778, 700, 779, 780, 781, 782]. This attack vector enables progressive erosion of defensive measures through iterative interaction patterns.
6.2.5 Defense.
@@ 6.2.6 Environment Safety
Perception The perception phase serves as the foundational layer of agent-environment interaction, where agents acquire raw data to interpret their surroundings. However, this phase is inherently vulnerable to risks such as data poisoning, environmental noise, and biased observations. Hudson [787] converts real-time sensory inputs into natural language representations augmented with security validation protocols, employing causal analysis techniques to improve reliability during adversarial perception scenarios. ChatScene [793] develops safety-oriented simulation environments for autonomous systems by converting linguistic commands into executable code compatible with CARLA’s simulation architecture. Chen et al. [800] systematically categorize perceptual vulnerabilities in financial AI systems, identifying three primary risk categories: synthetic data generation errors, temporal inconsistency challenges, and susceptibility to engineered input manipulations.
@@
Reasoning The reasoning phase transforms raw perceptual data into actionable insights through decision-making models, and knowledge-based inference. This stage is critical to ensure agents act appropriately in dynamic environments, but introduces unique trustworthiness challenges. Yang et al. [792] develop a temporal safety verification framework using formal logic systems, implementing dual mechanisms for auditing the compliance of safety protocols and filtration of hazardous decisions to meet the requirements of industrial robotics. Agents4PLC [801] establishes an industrial control programming framework that combines automated code synthesis with formal verification processes, integrating RAG [216] and COT [313] to ensure operational integrity. Xiang et al. [768] propose medical AI systems that employ semantic reasoning engines for confidential data protection. Park et al. [791] demonstrate improved threat detection capabilities through simulated organizational communication patterns in anomaly identification systems.
Action The action phase represents the culmination of agent-environment interaction, where agents execute decisions to influence their surroundings or update internal states. Trustworthiness at this stage hinges on ensuring that actions are safe, precise, and aligned with intended objectives. Fang et al. [797] reveal the capacity of autonomous systems to exploit digital infrastructure weaknesses through adaptive penetration testing, prompting the development of specialized evaluation frameworks for web agents. Furthermore, researchers develop frameworks to evaluate the truthfulness of web agents. Polaris [802] implements distributed AI architectures to enhance fault tolerance and response accuracy of healthcare interaction systems. La et al. [803] employ linguistic evolution models to simulate adaptive content generation patterns that circumvent automated moderation systems, providing insights for regulatory mechanism improvements.
@@ 6.3.1 Attack
Strategic Attack. Strategic attack involves collaboration between agents and strategic optimization of attack methods, aiming to emphasize the cooperation and long-term impact of the attack, making it increasingly dangerous and more destructive. Evil Geniuses [816] modifies system roles, where these roles collaborate to generate malicious prompts. By simulating adversarial attacks and defenses, they optimize and evaluate each round of attack behavior, making the attacks increasingly dangerous to target other agents. Amayuelas et al. [817] use adversarial attack techniques to enable harmful agents in the multi-agent system to collaborate in debates to persuade other secure agents. These malicious agents may exploit superior knowledge, larger model sizes, or greater persuasion power to gain an unfair advantage. Ju et al. [818] forms a multi-agent community using a two-stage attack method: persuasive injection and knowledge manipulation injection, to induce agents to spread counterfactual and harmful knowledge.
@@ 6.4 Agent Safety Evaluation
Currently, there is already a substantial body of work evaluating the performance of LLM-based agent systems on different tasks [827, 828, 829, 830]. In this section, we focus on benchmarks designed to assess the security of agents. Broadly speaking, these benchmarks include those that construct datasets and those that use other agents to set up sandbox environments for evaluation, each with distinct assessment priorities and specific scenarios for agent security [831, 832, 833, 285, 834].
@@
6.4.1 Attack-Specific Benchmarks.

### v2

@@ 6.1.2 Defensive Mechanisms in Deployment
Output Filtering Mechanisms. Output filtering mechanisms [212, 650] serve as a critical safeguard in LLM deployment, ensuring that generated responses comply with safety constraints while preserving informativeness. Unlike input preprocessing, which aims to prevent adversarial prompts from reaching the model, output filtering mitigates harmful content post-generation. Existing approaches primarily follow three paradigms: rule-based constraints, generative adversarial filtering, and toxicity detection.
@@
Toxicity detection [663, 664, 665] and content moderation [666, 667, 668, 669] further reinforce output safety by identifying and mitigating hate speech [670], misinformation, and other harmful content. Supervised fine-tuning adapts LLMs to recognize undesirable patterns, while classifier-based detection models [671] filter responses in real-time. Some approaches introduce debiasing strategies, such as controlled decoding [672, 673] and anti-expert guidance [674], to suppress toxic outputs without sacrificing response diversity. However, these methods face challenges in balancing false positives and false negatives, particularly in ambiguous or context-dependent cases.
@@
Robust Prompt Engineering. Robust prompt engineering aims to enhance LLM safety by designing input prompts that resist adversarial manipulation [675], protect sensitive data, and mitigate harmful outputs—all [676] without modifying model parameters. These strategies act at the interaction level, offering lightweight and model-agnostic protection.
@@
Privacy-preserving prompt [686] design introduces formal guarantees through differential privacy. Approaches like DP-Prompt [687] and stochastic gradient masking [688] reduce information leakage from prompts without harming performance. Desensitization and directional control of in-context representations offer additional privacy protections during prompt construction. Prompt engineering [572, 689] also helps mitigate societal risks. Chain-of-thought prompting and guided templates reduce gender bias [690] in reasoning tasks, while prompt learning [691] improves toxicity detection and generation control [692, 693], often surpassing specialized models in efficiency and generalization.
@@
System-level Security Controls. System-level defenses [698] enhance LLM deployment by optimizing inference, enforcing alignment, isolating untrusted inputs, and securing the supply chain. Systems like Petals [699], Sarathi-Serve [700], and DistServe [701] restructure computation to improve throughput and latency, while TriForce [702], Medusa [703] MagicDec [704] accelerate generation via speculative decoding and structural compression. Parallel frameworks such as DeepSpeed-FastGen [705] and SpecExec [706] further boost efficiency with minimal overhead.
@@
LLM-based guard models utilize lightweight LLMs like Llama Guard [323], Aegis Guard [712, 713], WildGuard [714], and ShieldGemma [715] to moderate both the input and output of the victim LLMs. However, they are purely classifiers. To solve this problem, the first reasoning-based guard model named GuardReasoner [716] is proposed to improve the performance, explainability, and generalization ability via learning to reason. It brings new opportunities for the safety of large-scale reasoning models [717].
@@ 6.1.3 Evaluation and Benchmarks in Deployment
To systematically assess the reliability of large language models (LLMs) after deployment, we categorize robustness evaluation into two broad types: adversarial robustness and natural robustness. Adversarial robustness focuses on evaluating how LLMs respond to malicious or adversarial inputs, such as jailbreak prompts, prompt injections, or red-teaming attacks. Natural robustness, on the other hand, assesses LLM behavior under non-malicious but realistic distribution shifts, including typos, paraphrasing, novel word usage, or temporal drift. A summary of representative benchmarks categorized along these 4 dimensions is presented in Table VII.
@@
Content Trustfulness and Fairness Evaluation.
@@
Data Privacy and Leakage Evaluation.
@@
Multi-modal Safety Evaluations As multimodal large language models (MLLMs) become increasingly integrated into real-world applications, ensuring their safety under diverse input conditions is essential. A growing number of studies have proposed evaluation benchmarks and frameworks to assess MLLM vulnerabilities across multiple dimensions [753, 754, 755, 756, 757, 758, 759, 760, 761, 762, 763, 764, 765, 766, 767, 768, 769, 770, 771, 772, 773, 774, 775].
@@ 6.2 Single-agent Safety
In this section, we focus on security issues related to a single agent. We first define an agent as an interactive entity that uses an LLM as the core for reasoning, decision-making, and reflection while integrating memory, tools, and the environment as capability-enhancing components. Beyond the deployment risks associated with the LLM core, we introduce the security issues arising from these three additional modules. Specifically, for tools (Section 6.2.2) and memory (Section 6.2.3), we summarize existing work from both attack (Section 6.2.4) and defense (Section 6.2.5) perspectives to identify technical paradigms. For the environment (Section 6.2.6), we explore unique security challenges from the perspective of various agent-interaction settings. We demonstrate an overview of agent safety in Figure 12.
@@ 6.2.2 Tool Safety
Attacks. Based on the target of the attack, safety-related attacks involving tools can be categorized into Tool-aided Attacks and Tool-targeted Attacks. The former refers to attackers utilizing agents equipped with tools to execute attacks that LLMs cannot independently assist with, such as leveraging agents with web access and code execution capabilities to facilitate cyberattacks. The latter involves attackers targeting the tool invocation process itself, attempting to manipulate or induce tool selection for malicious purposes through various attack methods. However, from the perspective of the technical stack of attacks, the two can be unified. We have identified new applications of traditional LLM attack methods in tool safety, as well as novel attack paradigms that have emerged due to the unique characteristics of tools.
@@
Defenses. Compared to attacks on agent tools, defense mechanisms for secure tool invocation have been less studied. Specifically, AgentGuard [813] employs LLM orchestrators to automatically detect unsafe tool-use workflows and produce safety constraints for secure tool utilization. PrivacyAsst [814] proposes an encryption-based solution by integrating an encryption scheme into the tool using LLM agents to safeguard user privacy and align them with computational security standards. In addition, some works enhance the security of agent systems by leveraging tool invocation, GuardAgent [815] pioneers an approach to verify target agents’ trustworthiness by executing guardrail code through API calls during task plan implementation.
@@ 6.2.4 Attack
(I) Memory Poisoning refers to adversarial attacks where malicious data is injected into an agent’s long-term memory [818, 819, 820, 821, 306, 822]. When the agent retrieves and utilizes such corrupted memory, it may produce erroneous outputs, misleading responses, or even hazardous actions. For example, PoisonedRAG framework [820] employs a dual optimization approach, simultaneously manipulating both the retrieval and generation pipelines to systematically poison the agent’s memory system. AgentPoison [819] introduces an advanced backdoor attack methodology that optimizes trigger patterns and seamlessly integrates them into query formulations, significantly elevating the likelihood of malicious sample retrieval while maintaining stealth. (II) Privacy Leakage occurs when attackers exploit the interface between an agent and its long-term memory to extract stored sensitive data [823, 600, 824, 513, 598]. Such breaches may expose user information to malicious third parties, posing significant real-world risks. (II) Memory Misuse refers to the deliberate construction of multi-turn query sequences that systematically circumvent safety protocols by exploiting the retention properties of agent short-term memory [825, 745, 826, 827, 828, 829]. This attack vector enables progressive erosion of defensive measures through iterative interaction patterns.
6.2.5 Defense
@@ 6.2.6 Environment Safety
Perception. The perception phase serves as the foundational layer of agent-environment interaction, where agents acquire raw data to interpret their surroundings. However, this phase is inherently vulnerable to risks such as data poisoning, environmental noise, and biased observations. Hudson [834] converts real-time sensory inputs into natural language representations augmented with security validation protocols, employing causal analysis techniques to improve reliability during adversarial perception scenarios. ChatScene [840] develops safety-oriented simulation environments for autonomous systems by converting linguistic commands into executable code compatible with CARLA’s simulation architecture. Chen et al. [847] systematically categorize perceptual vulnerabilities in financial AI systems, identifying three primary risk categories: synthetic data generation errors, temporal inconsistency challenges, and susceptibility to engineered input manipulations.
@@
Reasoning. The reasoning phase transforms raw perceptual data into actionable insights through decision-making models, and knowledge-based inference. This stage is critical to ensure agents act appropriately in dynamic environments, but introduces unique trustworthiness challenges. Yang et al. [839] develop a temporal safety verification framework using formal logic systems, implementing dual mechanisms for auditing the compliance of safety protocols and filtration of hazardous decisions to meet the requirements of industrial robotics. Agents4PLC [848] establishes an industrial control programming framework that combines automated code synthesis with formal verification processes, integrating RAG [235] and COT [336] to ensure operational integrity. Xiang et al. [815] propose medical AI systems that employ semantic reasoning engines for confidential data protection. Park et al. [838] demonstrate improved threat detection capabilities through simulated organizational communication patterns in anomaly identification systems.
Action. The action phase represents the culmination of agent-environment interaction, where agents execute decisions to influence their surroundings or update internal states. Trustworthiness at this stage hinges on ensuring that actions are safe, precise, and aligned with intended objectives. Fang et al. [844] reveal the capacity of autonomous systems to exploit digital infrastructure weaknesses through adaptive penetration testing, prompting the development of specialized evaluation frameworks for web agents. Furthermore, researchers develop frameworks to evaluate the truthfulness of web agents. Polaris [849] implements distributed AI architectures to enhance fault tolerance and response accuracy of healthcare interaction systems. La et al. [850] employ linguistic evolution models to simulate adaptive content generation patterns that circumvent automated moderation systems, providing insights for regulatory mechanism improvements.
@@ 6.3.1 Attack
Strategic Attack. Strategic attack involves collaboration between agents and strategic optimization of attack methods, aiming to emphasize the cooperation and long-term impact of the attack, making it increasingly dangerous and more destructive. Evil Geniuses [863] modifies system roles, where these roles collaborate to generate malicious prompts. By simulating adversarial attacks and defenses, they optimize and evaluate each round of attack behavior, making the attacks increasingly dangerous to target other agents. Amayuelas et al. [864] use adversarial attack techniques to enable harmful agents in the multi-agent system to collaborate in debates to persuade other secure agents. These malicious agents may exploit superior knowledge, larger model sizes, or greater persuasion power to gain an unfair advantage. Ju et al. [865] form a multi-agent community using a two-stage attack method: persuasive injection and knowledge manipulation injection, to induce agents to spread counterfactual and harmful knowledge.
@@ 6.3.2 Defense
6.4 Agent Communication Safety
As Large Language Model (LLM)-based Agents evolve from isolated entities into interconnected MAS, the mechanisms governing communication between Agents, and their interactions with external environments and tools, have become increasingly critical. Agents exchange information and collaborate through message passing, tool invocation, and environmental interactions; these mechanisms, while essential to system functionality, also expose significant attack surfaces. Early methods [874, 875, 876, 877, 878, 879, 880] of Agent interaction often relied on ad-hoc approaches, such as shared memory [881],API calls [882] or unstructured function calls [883], leading to fragmented systems lacking unified security considerations. To address this challenge and enhance interoperability, standardized communication protocols have emerged. Examples include Anthropic’s Model Context Protocol (MCP) [884] for Agent-tool interactions, Google’s Agent2Agent (A2A) [885] for enterprise-level Agent collaboration, and the Agent Network Protocol (ANP) [886] for open network interoperability, along with other commonly used protocols [887, 888, 889, 890, 891, 892, 893, 894, 895, 896, 897],However, the open design and dynamic nature of these communication mechanisms, coupled with the autonomy of the Agent, has exposed new vulnerabilities while enhancing functionality.
6.4.1 Attack
The interconnected nature of MAS, facilitated by numerous communication channels, creates a multifaceted attack surface. While individual Large Language Models (LLMs) possess inherent vulnerabilities, the interactions and communications among Agents introduce novel threats that exploit the system’s collaborative dynamics. These threats target various components, including communication channels, content interpretation, and underlying protocols, with examples such as Shadowing Attacks, Naming Attacks, Context Poisoning, and Rug Pulls.
Attacks Communication Channels. These attacks directly disrupt the transmission and routing of messages in the system, affecting both inter-Agent communications and interactions with external endpoints. For instance, Agent-in-the-Middle (AiTM) attacks [862] specifically target the core communication mechanisms of LLM-MAS. By intercepting and manipulating messages between Agents, these attacks can cause Agents to perform unintended actions, thereby compromising the entire system. Such attacks underscore the critical security vulnerabilities arising from the communication-dependent nature of Agent collaboration. Furthermore, attacks targeting communication channels and transmission processes, such as communication perturbation [898], involve adversaries injecting noise into messages in transit [899] or masquerading as legitimate sources [900], thereby compromising both the efficiency and security of Agent collaboration.
Attacks Content. These attacks target the content of messages themselves, leveraging the mechanisms by which Agents process and interpret received information. For example, Prompt Injection involves embedding malicious instructions into data or content that Agents retrieve or receive through communication channels, thereby manipulating the Agent’s behavior or decision-making processes. This technique is discussed in several works, such as [593] and [536]. Additionally, [901] explores indirect Prompt Injection within tool-based scenarios, highlighting the varied strategies employed in complex environments.
Attacks Exploiting Multi-Agent Dynamics. These attacks leverage the interconnected structure, interaction patterns, or collective behavior of communication-driven Multi-Agent Systems (MAS) to amplify their impact or achieve strategic objectives. Contagious attacks (propagation) initiate malicious behavior on a single agent and spread it across the entire network via inter-agent communication [858, 822]. Additionally, malicious agents can coordinate through collective communication to achieve harmful goals, such as replicating malicious instructions across the network by sending replication code or commands, thereby leading to the sharing of legitimate communication keys or identity information with other malicious entities [902].
6.4.2 Defense
To tackle threats to Agent communication, research proposes a multi-layered defense strategy addressing key points across the communication pipeline, from infrastructure to Agent-level processing. These defenses aim to prevent, detect, or mitigate attacks on channels, content, infrastructure, dynamics, and environmental factors. The strategies integrate into infrastructure and protocol design, individual Agents’ message processing, and the collaborative and learning mechanisms of the MAS.
Protocol Defenses. Protecting the foundation of Agent communication. This includes adopting standardized protocols with built-in security features (encryption, integrity checks, authentication) To counter Agent communication threats, research proposes multi-layered defense strategies targeting different points in the communication pipeline, from the underlying infrastructure to Agent-level message processing. Effective defenses aim to prevent, detect, or mitigate attacks on communication channels, content, infrastructure, such as MCP [884], A2A [885], ANP [886] standards. Establishing managed registries and identity systems for Agent and Tool/Service registration and identity management. Enforcing strong Agent identity verification and access control policies, including JIT credential provisioning. Implementing mechanisms to enforce communication dynamics, and environmental impacts.
Content Defense. These defenses operate at the agent level, focusing on how agents process received messages and content. This includes input modification and filtering, which preprocess incoming content to neutralize adversarial elements. Agents also employ active defense mechanisms, such as reliability estimation, to assess the trustworthiness of messages based on local context, thereby mitigating the impact of untrusted information. For example, [903] proposed an active defense strategy that utilizes a reliability estimator to judge the credibility of received messages and employs a decomposable message aggregation policy network to reduce the influence of unreliable messages on the final decision.
@@ 6.5 Agent Safety Evaluation
Currently, there is already a substantial body of work evaluating the performance of LLM-based agent systems on different tasks [904, 905, 906, 907, 908]. In this section, we focus on benchmarks designed to assess the security of agents. Broadly speaking, these benchmarks include those that construct datasets and those that use other agents to set up sandbox environments for evaluation, each with distinct assessment priorities and specific scenarios for agent security [909, 910, 911, 307, 912].
@@
6.5.1 Attack-Specific Benchmarks